OpenStudy (happykiddo):

A square pyramid of height 8 inches has a volume of 704 cubic inches. If the volume of a similar square pyramid is 88 cubic inches, what is its height? 1 inch 2 inches 4 inches 8 inches

OpenStudy (accessdenied):

What would be the ratio of their volumes?

OpenStudy (happykiddo):

oh gosh i forgot about using ratios thanks

OpenStudy (accessdenied):

If we know the ratio of the volumes, we can find the ratio of the dimensions by taking the cube root. \( {\text{Ratio of volumes:} \frac{a^3}{b^3}\\ \implies \text{Ratio of dimensions:} \frac{a}{b}} \) Then just use a proportion to find the height of the smaller pyramid: \( \text{Ratio of dimensions} = \frac{8}{h} \)

OpenStudy (accessdenied):

Oh, okay. You're welcome! :P